OpenStudy (anonymous):

!!!!GEOMETRY HELP ASAP!!!! Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 − 4x + y2 + 8y = −4 I dont understand how to do this))):

OpenStudy (anonymous):

Do you know the equation for the center of a circle?

OpenStudy (anonymous):

um is it (x-h)2+(y-k)2

OpenStudy (anonymous):

You have to complete squares: (a + b)^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

So what do I do??

OpenStudy (anonymous):

x^2 − 4x + y^2 + 8y = −4, completing the squares: (x - 2)^2 = x^2 -4x + 4 (y + 4)^2 = y^2 + 8y + 16, so... x^2 − 4x + y^2 + 8y = −4 (x-2)^2 - 4 + (y+4)^2 - 16 = -4 (x-2)^2 + (y+4)^2 = 16 (x-2)^2 + (y+4)^2 = 4^2 that means: Center of circle = (2, -4) Radius = 4

OpenStudy (anonymous):

tht ll ame out as like question mark symbols...

OpenStudy (anonymous):